How much can you borrow if you are willing to make a single payment of 100,000 twelve years from now and the lender will charge 6% annual interest, compounded quarterly? what is the total amount of interest you will pay
Answer #1

Interest rate Annually = 6%

Therefore for Quarterly = 4 months / 12 months * 6% = 1.5%

Number of times Interest is to be charged = 12 years * 4 Quarters = 48 quarters

Amount Borrowed today = = 0.48936 *100000 = 48936

Excel Formula = (1/(1+(1*1.5/100))^48)*100000

Total Interest to be paid = 100000-48936 = 51064

Following is a table for reference

No of Quarters Principle Interest Cumulative
1        48,936            734            49,670
2        49,670            745            50,415
3        50,415            756            51,171
4        51,171            768            51,939
5        51,939            779            52,718
6        52,718            791            53,509
7        53,509            803            54,311
8        54,311            815            55,126
9        55,126            827            55,953
10        55,953            839            56,792
11        56,792            852            57,644
12        57,644            865            58,509
13        58,509            878            59,386
14        59,386            891            60,277
15        60,277            904            61,181
16        61,181            918            62,099
17        62,099            931            63,031
18        63,031            945            63,976
19        63,976            960            64,936
20        64,936            974            65,910
21        65,910            989            66,898
22        66,898        1,003            67,902
23        67,902        1,019            68,920
24        68,920        1,034            69,954
25        69,954        1,049            71,003
26        71,003        1,065            72,069
27        72,069        1,081            73,150
28        73,150        1,097            74,247
29        74,247        1,114            75,360
30        75,360        1,130            76,491
31        76,491        1,147            77,638
32        77,638        1,165            78,803
33        78,803        1,182            79,985
34        79,985        1,200            81,185
35        81,185        1,218            82,402
36        82,402        1,236            83,638
37        83,638        1,255            84,893
38        84,893        1,273            86,166
39        86,166        1,292            87,459
40        87,459        1,312            88,771
41        88,771        1,332            90,102
42        90,102        1,352            91,454
43        91,454        1,372            92,826
44        92,826        1,392            94,218
45        94,218        1,413            95,631
46        95,631        1,434            97,066
47        97,066        1,456            98,522
48        98,522        1,478        1,00,000